# Signal Processing over Time-Varying Graphs: A Systematic Review

### v1

@@ V-B Dynamic Graph Neural Networks
2) CTDG Methods: Since the graph representation is totally different from DTDG, where the event set stores the temporal information rather than iteratively forwarding previous information like RNN. As a consequence, the DGNN methods for CTDG are more diverse. A class of methods models the event sequence by temporal point process, where the occurrence of events recording the evolving graph signal is assumed to follow a multivariate temporal point process. The conditional intensity function of this process can be modulated by the score of the relevant relationship, and the measurement of the relationship can be realized by GNNs [117, 118]. Besides the temporal point process, as an extension of random walk [26, 27], the temporal random walk is utilized for learning spatiotemporal association as well. In temporal random walk, the movement from one vertex to another will depend on its spatiotemporal neighbors to extract dynamics from the time-varying graph [119, 120].
@@
On the other hand, some researchers also model the relationship between temporal events directly by neural networks. For instance, the Dynamic Graph Neural Networks (DyGNN) [121] employ an update component and a propagation component to cope with coming events. As shown in Fig. 5(c), when a new event comes, the update component maintains the freshness of node information by learning the relationship with previously memorized events and the time interval between the stored events and the new event. Then, the propagation component propagates the updated information to the affected nodes (e.g., its neighborhood nodes) by considering the intensity of the influence. Meanwhile, JODIE [122] is developed and shares a similar scheme with DyGNN, which also has the update operation and the projection operation. The major difference is, in JODIE, each node has a static embedding to represent its fixed attributes, and a dynamic embedding to reflect its current state.
Remarks: DGNN is a sort of more flexible NN method for learning and processing time-varying graphs compared with STGNNs, by considering the potential temporal variation in high-order structures of time-varying graphs. However, more complicated modeling generally leads to higher complexity. For instance, in the traffic network, the topology is hard to change, whereas STGNNs are more popular owing to their efficiency. Thus, selecting a proper class of models according to the graph data encountered in practice is critical.
V-C Graph Lifelong Learning
@@
Graph lifelong learning [123] (a.k.a., continual graph learning [124]) is an emerging research topic and mainly developed for processing real-world time-growing graphs. The scene of graph lifelong learning is that the number of nodes is growing over time, or there will be new nodes of unseen classes. As depicted in Fig. 5(c), it solves this problem by iteratively training the GNN. Like general continual learning, the core problem in graph lifelong learning is also catastrophic forgetting. To prevent this, the current graph lifelong learning methods mostly refer to continual learning and can be classified into three categories including the architectural methods, the regularization methods, and the experience-replay methods.
@@
1) Architectural Methods: The architectural methods [123, 125] aim to change the architecture of the GNN iteratively to coherently adapt to the changes that occurred in streaming data. The Feature Graph Networks (FGN) [123] are a special type of architectural method to tackle time-growing graphs by node-feature transformation. Specifically, it first transforms each node into a feature graph, and the features in this node become new nodes in each feature graph. The edge weights between these feature nodes are determined by cross-correlation. Specifically, if the feature node matrix is , then the feature adjacency matrix can be defined as:
@@
where represents the k-hop neighbor of , indexes the channel. In this way, the continually coming nodes can be regarded as an individual graph and classified with previous techniques in continual learning.
@@
2) Experience-Replay Methods: In addition to architectural methods, like in continual learning, the experience replay technique is also employed for graph continual learning [126, 127]. In the Experience Replay Graph Neural Networks (ER-GNN) [126], a set of nodes are stored in a buffer as experience and replayed in the subsequent task to avoid forgetting. To identify valuable nodes of each task, ER-GNN proposes several metrics including the mean of features, coverage maximization, and influence maximization. Once the nodes are selected and stored in a buffer, the training objective for each task is altered as:
@@
where indicates the loss function (e.g., cross entropy in node classification tasks); indexes the task; represents the experience buffer; denotes the graph neural network; represents the training dataset of task .
@@
3) Regularization Methods: Adding regularization terms to penalize dramatic updates on important parameters is also common in continual learning [124, 128, 129]. The representative one in this line of work is, a plug-and-play module, called “Topology-aware Weight Preserving” (TWP) [124], which estimates an importance score for each network parameter. This score is computed as the sum of gradients concerning parameters (denoted as ) and gradients concerning attention coefficients parameters (denoted as ). The first term aims to minimize loss of different tasks, and it can be calculated from:
@@
where represents a single neural network layer for feature projection; represents the concatenated feature embedding of node and at -th layer in the graph attention neural network.
@@
Remarks: Dealing with the time-varying graph by graph continual learning is a noteworthy direction, which merges techniques from both continual learning and graph neural networks and achieves great performance in some specific scenarios (e.g., learning and processing the time-growing graphs [123]).
| Type | Dataset | Domain | Catagory | #Nodes | #Edges | #Timestamps |

### v2

@@ 4.2.2 CTDG Methods
Since the graph representation is different from DTDG, where the event set stores the temporal information rather than iteratively forwarding previous information like RNN. As a consequence, the DGNN methods for CTDG are more diverse. A class of methods models the event sequence by a temporal point process, where the occurrence of events recording the evolving graph signal is assumed to follow a multivariate temporal point process. The conditional intensity function of this process can be modulated by the score of the relevant relationship, and the measurement of the relationship can be realized by GNNs [179, 180]. Besides the temporal point process, as an extension of random walk [136, 55], the temporal random walk is utilized for learning spatiotemporal association as well. In a temporal random walk, the movement from one vertex to another will depend on its spatiotemporal neighbors to extract dynamics from the TVG [190, 104].
@@
On the other hand, some researchers also model the relationship between temporal events directly by neural networks. For instance, the Dynamic Graph Neural Networks (DyGNN) [112] employ an update component and a propagation component to cope with coming events. As shown in (c) of Fig. 2, when a new event comes, the update component maintains the freshness of node information by learning the relationship with previously memorized events and the time interval between the stored events and the new event. Then, the propagation component propagates the updated information to the affected nodes (e.g., its neighborhood nodes) by considering the intensity of the influence. Meanwhile, JODIE [90] is developed and shares a similar scheme with DyGNN, which also has the update operation and the projection operation. The major difference is that in JODIE, each node has a static embedding to represent its fixed attributes and a dynamic embedding to reflect its current state.
4.3 Graph Lifelong Learning
@@
Graph lifelong learning [185] (a.k.a., continual graph learning [221]) is an emerging research topic and is mainly developed for processing real-world time-growing graphs. The scene of graph lifelong learning is that the number of nodes is growing over time, or there will be new nodes of unseen classes. As depicted in (d) of Fig. 2, it solves this problem by iteratively training the GNN. Like general continual learning, the core problem in graph lifelong learning is also catastrophic forgetting. To prevent this, the current graph lifelong learning methods mostly refers to continual learning and can be classified into three categories, including the architectural methods, the regularization methods, and the experience-replay methods.
4.3.1 Architectural Methods
@@
The architectural methods [185, 174] aim to change the architecture of the GNN iteratively to coherently adapt to the dynamic data stream. The Feature Graph Network [185] is a special type of architectural method to tackle time-growing graphs by node feature transformation. Specifically, it transforms each node into a feature graph, and the features in this node become new nodes in each feature graph. The edge weights between these feature nodes are determined by cross-correlation. Specifically, if the feature node matrix is , the feature adjacency matrix is
@@
where represents the -hop neighbor of , indexes the channel, and is the number of -hop neighbor of . In this way, the continually coming nodes are regarded as an individual graph and classified with previous techniques in continual learning.
4.3.2 Experience-Replay Methods
@@
In addition to architectural methods, like in continual learning, the experience replay technique is also employed for graph continual learning [106, 45]. In the Experience Replay Graph Neural Networks (ER-GNN) [106], a set of nodes is stored in a buffer as experience and replayed in the subsequent task to avoid forgetting. To identify valuable nodes of each task, ER-GNN proposes several metrics, including the mean of features, coverage maximization, and influence maximization. Once the nodes are selected and stored in a buffer, the training objective for each task is altered as:
@@
where indicates the loss function (e.g., cross entropy in node classification tasks); indexes the task; represents the experience buffer; denotes the GNN; represents the training dataset of task .
@@
Regularization Methods Adding regularization terms to penalize dramatic updates on important parameters is common in continual learning [221, 175, 87]. The representative one in this line of work is a plug-and-play module called Topology-aware Weight Preserving [221], which estimates an importance score for each network parameter. This score is computed as the sum of gradients concerning parameters (denoted as ) and gradients concerning attention coefficients parameters (denoted as ). The first term aims to minimize the loss of different tasks, and it can be calculated from:
@@
where represents a single neural network layer for feature projection, and are gradient losses, represents the concatenated feature embedding of node and at layer in the graph attention neural network. Dealing with the TVG by graph continual learning is a noteworthy direction, which merges techniques from both continual learning and GNNs and achieves great performance in some specific scenarios (e.g., learning and processing the time-growing graphs [185]).
5 Applications
On the application side, recent advancements in TVGSP and TVGNNs have driven progress in various fields, including social science, biological analysis, transportation, environmental studies, and finance. This growth is attributed to the exceptional capability of TVG representation in capturing dynamic relationships among irregularly distributed signals [157, 167, 130, 39]. Here, we will be providing a discussion on how TVGSP and TVGNN algorithms are utilized.
5.1 Social Science
In social science, TVGSP and TVGNN emerged as powerful tools for analyzing dynamic datasets, offering insights into evolving relationships, behaviors, and social interaction patterns. Applications in this domain span various areas, such as opinion dynamics, community detection, anomaly detection, and modeling social behavior.
5.1.1 Opinion Dynamics and Influence Propagation:
Opinion dynamics and influence propagation in social networks can be modeled using TVGs, where nodes represent users and edges represent temporal interactions. TVGs enable tracking the spread and evolution of opinions or sentiments. Dynamic graph embeddings on platforms like Reddit and Wikipedia showcase this [165, 91].
5.1.2 Community Detection:
TVGSP enables the monitoring of changes, mergers, or splits of communities in social networks. Community detection in dynamic networks benefits greatly from TVG-based methodologies, which allow for the identification and tracking of evolving communities over time in physical and digital contexts. For example, TVG analysis enables researchers to detect emerging communities and shifting user preferences over time, providing valuable insight into online social interactions [132, 158].
5.1.3 Detection of Anomalous Patterns:
TVGSP is highly useful for anomaly detection in social networks, where unusual patterns in user behavior or interactions can be flagged in real time. Such capabilities are crucial for monitoring social media for events like coordinated misinformation efforts or network breaches. In the financial sectors, the DGraphFin [71] dataset supports such TVGSP applications, where anomaly detection algorithms are used to detect fraud and irregular transaction patterns. In social contexts, [10] introduces a centrality measure designed for continuously updating network streams, helping identify influential nodes and interactions as new data arrives. [14] employed a multi-level anomaly detection algorithm in dynamic graph settings, effectively detecting anomalies at different graph levels (e.g., node, subgraph) with interactive visualization. These methods prove valuable for anomaly detection within evolving social graphs where central actors can be continually monitored.
5.2 Biomedical analysis
Dynamic graph-based methods are particularly suited to represent temporally evolving interactions, such as neural connectivity patterns, gene expression, and physiological signals. Biomedical applications of TVGSP encompass brain connectivity studies, predictive modeling in medical diagnostics, and protein interaction analysis.
5.2.1 Dynamic Functional Brain Connectivity Analysis:
Functional connectivity analysis is an increasingly important method for studying brain function, examining the statistical codependencies between signals from different brain regions [64]. TVG has become a valuable representation for analyzing functional brain connectivity, particularly in studies using time-series data from fMRI or EEG. Yu et al. [209] used TVGs to assess dynamic brain connectivity in patients with schizophrenia by analyzing time-varying functional networks derived from fMRI data, highlighting how TVG-based approaches identify connectivity alterations associated with mental disorders. Calhoun and Adali further expanded by using TVG-based techniques to track connectivity states across different brain regions, demonstrating applications in both healthy and clinical populations [18]. Zhao et al. achieved superior performance in brain disorder classification by integrating a particle filtering algorithm with GNNs [215]. Hou et al. integrated the generation, clustering and classification into one framework by learning the spectral dynamics and optimal clustering [65]
5.2.2 Medical Diagnostics Modeling:
In medical diagnostics, TVGSP and TVGNNs have enabled predictive modeling using dynamic patient data, offering tools for early diagnosis and outcome prediction. Xie et al. [194] introduced a graph convolutional neural network that predicts intensive care interventions based on temporal multivariate data from ICU records, demonstrating the potential for real-time decision support in critical care environments. Li et al. [62] explored TVGs in EEG analysis to monitor cognitive health and diagnose neurological disorders, showing that dynamic graphs can reveal early-stage signs of brain dysfunction in clinical populations.
5.2.3 Protein Interaction and Structural Dynamics Analysis:
TVGs provided a robust approach for studying dynamic protein interactions, enabling the modeling of temporal and structural changes within protein-protein interaction (PPI) networks. By applying TVGSP and TVGNNs, researchers can capture the evolution of interactions among proteins, which is essential for understanding functional pathways and responses to various stimuli. Song et al. [169] applied dynamic Bayesian networks to track protein signaling pathways over time, using TVGs to model interactions during cellular processes. Li et al. [99] demonstrated how TVGs can track dynamic changes in PPI networks, highlighting structural adaptations in proteins under different environmental conditions. Calazans et al. [17] leveraged TVGSP and TVGNNs for protein interaction analysis, using graph learning techniques to identify structural motifs and interaction dynamics in protein complexes.
5.3 Transportation
In transportation systems, TVGs have become an essential tool for analyzing and capturing the complex and evolving relationships between various components, such as vehicles, traffic signals, and road networks. Sensors along the roads are often modeled as graph nodes or edges, with the captured sensor readings treated as time-varying signals. Key applications include traffic flow prediction, congestion management, and route optimization.
5.3.1 Traffic Flow Prediction and Congestion Analysis:
Traffic flow prediction is one of the most prominent applications of TVGSP and TVGNNs. By treating traffic flow as a signal on a graph, TVG-based models capture both spatial dependencies (interconnectedness of adjacent road segments) and temporal variations (changes in flow over time). The widely used PeMS datasets from California, such as PeMSD7 [207] and PeMSBay [103], contain rich traffic sensor data, enabling effective applications of GSP for short- and long-term traffic prediction. For example, Li et al. [103] utilized GCNs on the PeMS datasets to accurately forecast traffic flows by learning spatial-temporal dependencies across the network.
5.3.2 Transportation Network Analysis:
TVGSP and TVGNN methods support the analysis and optimization of public transportation networks, where interactions between transit nodes (e.g., bus stops, train stations) and connections (routes) change continuously over time [148]. By analyzing passenger demand patterns, TVG-based methods can optimize transit scheduling and enhance network efficiency. For instance, GPS-based movement data from taxis are analyzed by TVGCNs to extract passenger demand patterns and recommend efficient resource allocation strategies [216]. Besides, TVGSP algorithms have also been applied for the analysis of GPS displacement [138]
5.3.3 Route Optimization and Navigation:
For real-time route optimization, TVG approaches with graphs that have edge weights that vary with traffic conditions, road closures, or accidents, enabling route recommendations and adapting to evolving traffic patterns. For instance, STGs are used to model real-time traffic speed data from Los Angeles, enabling adaptive routing recommendations based on evolving traffic patterns and optimizing travel routes in real time [103]. TVGs provide predictive frameworks for identifying probable path failure, supporting route planning that can proactively avoid high-risk paths [97].
5.4 Environmental Science
Evolving interactions of the sensors on sensor grids across different locations naturally form a graph based on sensor locations. TVGs serve as valuable frameworks for monitoring and modeling natural processes, with applications in air quality forecasting, Meteorological prediction, and environmental hazard detection.
5.4.1 Air Quality Prediction:
TVG approaches enable the analysis of air pollution signals, providing real-time insights into pollutant levels. A model for reconstructing TVG signals was introduced by Qiu et al. [140], which was applied to PM2.5 pollution data from California. This model enables enhanced tracking of pollution levels over time by leveraging the spatial dependencies between monitoring stations. A Sobolev smoothness-based framework demonstrated effective signal recovery in environmental datasets, highlighting its application to air quality data [50]. Spatiotemporal smoothness facilitates predictions of air quality by modeling correlations across sensor networks and adapting to temporal changes [108].
5.4.2 Weather Forecasting:
TVGs play a crucial role in analyzing Meteorological data. Castro-Correa et al. [19] advanced this area by developing Gegenbauer GNNs specifically designed for reconstructing time-varying signals in complex climate data, allowing for the extraction of meaningful trends over extended periods. There are various meteorological prediction scenarios, such as wind prediction [199, 63, 172, 142] and temperature prediction [170, 78, 22, 197], illustrating the capability of TVGSP in practice.
5.4.3 Environmental Hazard Detection:
Monitoring environmental hazards, such as floods, wildfires, and industrial pollution, benefits from the dynamic adaptability of TVGSP frameworks, which allow for the real-time detection of anomalies and the tracking of evolving hazard patterns. Valdivia et al. [181] utilized wavelet-based methods to visualize and process time-varying environmental data, enabling real-time analysis of environmental events such as wildfires. In a more recent study, Sun et al. [173] developed the Time-Varying Graph Convolutional Network (TVGCN) model to predict multivariate spatiotemporal series, which is particularly useful in forecasting hazard-related data such as rainfall and water levels in flood-prone areas.
5.5 Financial Analysis:
The integration of TVGs within financial analysis addresses diverse and complex challenges within dynamic financial systems, where interactions between entities such as stocks, transactions, and institutions change over time. TVGs empower researchers and analysts to harness the temporal and structural dynamics of financial interactions, enabling real-time insights into evolving networks. Financial applications of TVGSP and TVGNNs include fraud detection, market volatility analysis, and stock price forecasting.
5.5.1 Fraud and Anomaly Detection:
Financial networks are extensively used for identifying irregular transactions and preventing fraud [211]. Kurman et al. provided a framework for detecting fraudulent users through temporal analysis of user ratings, focusing on anomalies in user behaviors and their interactions within the network [88]. Weber et al. used GNNs on a temporal dataset of Bitcoin transactions to detect and prevent money laundering activities in digital finance, leveraging the Bitcoin and Elliptic dataset to identify laundering activities [191].
5.5.2 Market Volatility Analysis and Risk Assessment:
Predicting stock market volatility often requires monitoring dynamic, interdependent relationships across different stocks or sectors. The temporal dynamics captured through TVGs provide insights into correlations and dependencies that influence market risk and volatility. For instance, a TVG model can be employed to analyze the dynamic correlations in global stock indices, enabling enhanced forecasting capabilities for market shifts and volatility [24]. These edge weight prediction techniques help to understand market volatility by focusing on how relationship strengths evolve, offering predictive insights into potential price fluctuations. The incorporation of topological data analysis tools further elaborates on how changes in the topology and geometry of financial market correlations, particularly during a crash, can be crucial for anticipating drastic market changes and predicting black swan events [206]. Figure 7 illustrates the correlation among selected CSI 300 stocks, which exemplifies the dynamic correlations that can be analyzed through TVGs.
5.5.3 Stock Price Prediction:
Stock price prediction is a crucial component of financial trading, offering traders valuable insights for making informed decisions on buying, selling, or holding stocks. Accurate forecasts are critical in optimizing trading strategies and maximizing profitability. A common approach to building the TVG models for stock price prediction is to combine the GNNs (e.g., GCN, GAT) with time series predicting methods (e.g., RNN, LSTM) [105, 219].
6 Datasets
With growing interest in dynamic graphs, various time-varying datasets have emerged for evaluating model performance. We classify them into three types, STGs, DTDGs, and CTDGs. Here we briefly review the datasets; a summary is in Table 1.
6.1 Datasets for STG:
As can be seen from Table 1, most of the STG datasets belong to traffic data, with a few datasets for epidemic and neuroimaging. METR-LA [103] contains data collected from sensors installed on highways in Los Angeles County, capturing traffic speed at various locations at 5-minute intervals. Seattle Loop [30, 29] provides 5-minute interval spatiotemporal speed data collected from inductive loop detectors on Seattle freeways. It features averaged speed measurements from multiple detectors on main lanes in the same direction at specific mileposts. See Fig. 6 for an illustration. Chickenpox [149] is a spatiotemporal dataset of weekly chickenpox (childhood disease) cases from Hungary. The dataset consists of a county-level adjacency matrix and a time series of the county-level reported cases between 2005 and 2015. HCP [182] (Human Connectome Project) is a comprehensive collection of neuroimaging data designed to map the structural and functional connectivity of the human brain from over 1,200 healthy adult participants.
6.2 Datasets for CTDG:
The datasets in CTDG cover a wide range of domain types, such as proximity, finance, social, and so on. We chose some representative datasets for each domain to introduce them. Social Evolution [97] records dynamic interactions among students at MIT’s Human Dynamics Lab. Collected via mobile devices, it captures temporal social interactions such as phone calls, text messages, and physical proximity over several months. Bitcoin-Alpha [89, 88] is a dynamic network of trust relationships between users on the Bitcoin Alpha platform, where users rate one another based on the trustworthiness of cryptocurrency transactions. LastFM [91] contains one month of who-listens-to-which song information, which was timestamped interactions between users and artists on the LastFM music streaming platform, capturing user preferences over time. UCI [132] is an online temporal communication, such as emails, between students from the University of California, Irvine, with edges in the graph corresponding to messages exchanged between users. Taobao [81, 224] collected from the Chinese e-commerce platform Taobao, contains temporal interactions between users and items, representing user behavior in an online marketplace, which links the meaning of the type of action (e.g., clicking, purchasing, or adding to a wishlist).
6.3 Datasets for DTDG:
In the field of DTDG, the following datasets have become widely used, allowing researchers to analyze the evolution of relationships, dependencies, and interactions over discrete time intervals. UN Vote [4] captures voting records of countries in the United Nations General Assembly from 1946 onwards, with a node representing a country and an edge between nodes reflecting similar voting behavior in a given year. Elliptic [191] contains Bitcoin transaction data from the Elliptic platform, with each transaction labeled as either "licit"(e.g., exchanges, wallet providers, miners) or "illicit"(e.g., scams, malware, ransomware). Nodes represent Bitcoin addresses, and edges capture payment flow between these addresses in discrete time steps. Twitter-Tennis [10] contains interactions on Twitter related to tennis events, where nodes represent Twitter users and edges represent mentions, replies, or retweets within specific time intervals. ABIDE [37, 215] (Autism Brain Imaging Data Exchange dataset) consists of functional MRI (fMRI) data from individuals diagnosed with autism spectrum disorder and neurotypical controls, which is often used to create brain connectivity networks over time, where nodes represent brain regions and edges represent functional connectivity. As shown in Fig. 8, the illustration presents COVID-19 cases at the nodes and population mobility along the edges at various timestamps [201].
| Type | Dataset | Domain | Category | #Nodes | #Edges | #Timestamps |
@@ 6.3 Datasets for DTDG:
| Seattle Loop [30, 29] | Traffic | general | 323 | 762 | 26,304 | |
| PedalMe [148] | Traffic | general | 15 | 225 | * | |
| US Weather [131] | Meteorological | general | * | * | * | |
| Chickenpox [149] | Epidemiology | general | 20 | 61 | 522 | |